# Randomly punctured Reed–Solomon codes achieve list-decoding capacity over linear-sized fields

### v5

@@ 2.1 Basic notation
For positive integers $t$, let $[t]$ denote the set $\{1,2,\dots,t\}$. The Hamming distance $d(x,y)$ between two vectors $x,y\in\mathbb{F}_{q}^{n}$ is the number of indices $i$ where $x_{i}\neq y_{i}$. For a finite field $\mathbb{F}_{q}$, we follow the standard notation that $\mathbb{F}_{q}[X_{1},\dots,X_{n}]$ denotes the ring of multivariate polynomials with variables $X_{1},\dots,X_{n}$ over $\mathbb{F}_{q}$, and $\mathbb{F}_{q}(X_{1},\dots,X_{n})$ denotes the field of fractions of the polynomial ring $\mathbb{F}_{q}[X_{1},\dots,X_{n}]$. By abuse of notation, we let $X_{\leq i}$ or $X_{[i]}$ to denote the sequence $X_{1},\dots,X_{i}$, and we let, for example, $X_{\leq i}=\alpha_{\leq i}$ to denote $X_{1}=\alpha_{1},X_{2}=\alpha_{2},\dots,X_{i}=\alpha_{i}$. Given a matrix $M$ over the field of fractions $\mathbb{F}_{q}(X_{1},\dots,X_{n})$ and field elements $\alpha_{1},\dots,\alpha_{i}\in\mathbb{F}_{q}$, let $M(X_{\leq i}=\alpha_{\leq i})$ denote the matrix over $\mathbb{F}_{q}(X_{i+1},X_{i+2},\dots,X_{n})$ obtained by setting $X_{\leq i}=\alpha_{\leq i}$ in $M$.
@@ 2.2 Hypergraphs and connectivity
In this work, we maintain a hypergraph perspective of the list-decoding problem, which was introduced in [${}^{+}$22]. We describe a bad list-decoding instance with a hypergraph where the $L+1$ bad codewords identify the vertices and the $n$ evaluation points identify the hyperedges (Definition 2.1). While prior works described a bad list-decoding instance by $L+1$ sets indicating the agreements of the codewords with the received word, this hypergraph perspective gives us several advantages:
@@
    The constraints imposed by a bad list-decoding configuration yield a hypergraph that is weakly-partition-connected. This is a natural notion of hypergraph connectivity, which is well-studied in combinatorics [FKK03b, FKK03a, Kir03] and optimization [JMS03, FK09, Fra11, CX18], and which generalizes a well-known notion ($k$-partition-connectivity) for graphs [NW61, Tut61]. [^5] The notion of weakly-partition-connected sits between two other well-studied notions: $k$-partition-connected implies $k$-weakly-partition-connected implies $k$-edge-connected [Kir03]. Each of these three notions generalizes an analogous notion on graphs. On graphs, $k$-partition-connected and $k$-weakly-partition-connected are equivalent. This connection allows us to have more compact notation, definitions, and lemma statements.
@@
    Because we work with weakly-partition-connected hypergraphs, we save a factor of $L$ in Lemma 2.14 compared to the analogous lemma in [GZ23]. This allows us to improve the constant factor in alphabet size for Reed–Solomon codes from $2^{O(L^{2}/\varepsilon)}$ in [GZ23] to $2^{O(L/\varepsilon)}$ in Theorem 1.1.
  3. 3.
    For similar reasons, for random linear codes, the hypergraph perspective saves a factor of $L$ in the alphabet size exponent, improving from $2^{O(L^{2}/\varepsilon)}$ to $2^{O(L/\varepsilon)}$ in Theorem 1.3.
  4. 4.
    With the hypergraph perspective, we can give a new presentation of the results in [BGM23] and more directly connect the list-decoding problem to the GM-MDS theorem [DSY14, Lov18, YH19], as the heavy-lifting in the combinatorics is done using known results on hypergraph orientations. This is done in Appendix A.
A hypergraph $\mathcal{H}=(V,\mathcal{E})$ is given by a set of vertices $V$ and a set $\mathcal{E}$ of (hyper)edges, which are subsets of the vertices $V$. In this work, all hypergraphs have labeled edges, meaning we enumerate our edges $e_{i}$ by distinct indices $i$ from some set, typically $[n]$, in which case we may also think of $\mathcal{E}$ as a tuple $(e_{1},\dots,e_{n})$. Throughout this paper, the vertex set $V$ is typically $[t]$ for some positive integer $t$. The weight of a hyperedge $e$ is $\operatorname*{wt}(e)\stackrel{{\scriptstyle\rm def}}{{=}}\max(0,|e|-1)$, and the weight of a set of hyperedges $\mathcal{E}$ is simply $\operatorname*{wt}(\mathcal{E})\stackrel{{\scriptstyle\rm def}}{{=}}\sum_{e\in\mathcal{E}}{\operatorname*{wt}(e)}$.
[…]
@@
A key property of hypergraphs that we are concerned with is weak-partition-connectivity.
Definition 2.2 (Weak Partition Connectivity).
@@
where $|\mathcal{P}|$ is the number of parts of the partition, and $|\mathcal{P}(e)|$ is the number of parts of the partition that edge $e$ intersects.
To give some intuition for weak partition connectivity, we state two of its combinatorial implications. First, if a graph is $k$-weakly-partition-connected, then it is $k$-edge-connected [Kir03], which, by the Hypergraph Menger’s (Max-Flow-Min-Cut) theorem [Kir03, Theorem 1.11], equivalently means that every pair of vertices has $k$ edge-disjoint (hyper)paths between them. [^6] In general the converse is not true. Second, suppose we replace every hyperedge $e$ with an arbitrary spanning tree of its vertices (which we effectively do in Definition 2.6). The resulting (non-hyper)graph is $k$-partition-connected, [^7] In (non-hyper)graphs, $k$-partition-connectivity and $k$-weak-partition-connectivity are equivalent. which, by the Nash-Williams-Tutte Tree-Packing theorem [NW61, Tut61], equivalently means there are $k$ edge-disjoint spanning trees (this connection was used in [${}^{+}$22]).
@@
Lemma 2.3 (Bad list gives $k$-weakly-partition-connected hypergraph. See also Lemma 7.4 of [${}^{+}$22]).
@@
Lemma 2.3 follows from the following result about weakly-particion-connected hypergraphs
@@
Let $\mathcal{H}=(V,\mathcal{E})$ be a hypergraph with at least two vertices and with total edge weight $\sum_{e\in\mathcal{E}}\operatorname*{wt}(e)\geq k\cdot(|V|-1)$, for some positive integer $k$. Then there exists a subset $V^{\prime}\subset V$ of at least two vertices such that the hypergraph $\mathcal{H}^{\prime}=(V^{\prime},\{V^{\prime}\cap e:e\in\mathcal{E}\})$ is $k$-weakly-partition-connected.
@@
By assumption, $V^{\prime}=[L+1]$ satisfies (3), so $V^{\prime}$ exists (note that singleton subsets of $[L+1]$ satisfy (3) with equality). Let $\mathcal{H}=(V^{\prime},\mathcal{E}^{\prime})$ be the hypergraph with edge set =$\mathcal{E}^{\prime}=\{V^{\prime}\cap e:e\in\mathcal{E}\}$. By minimality of $V^{\prime}$, for all $V^{\prime\prime}\subsetneq V^{\prime}$, we have $\sum_{e\in\mathcal{E}^{\prime}}{\operatorname*{wt}(e\cap V^{\prime\prime})}\leq(|V^{\prime\prime}|-1)k$. Now, consider a non-trivial partition $\mathcal{P}=P_{1}\sqcup\cdots\sqcup P_{p}$ of $V^{\prime}$ where $P_{i}\neq V^{\prime}$ for all $i\in[p]$ (as otherwise (2) trivially follows). We have
$\sum_{e\in\mathcal{E}^{\prime}}{\max\{|\mathcal{P}(e)|-1,0\}}$ $=\sum_{e\in\mathcal{E}^{\prime},e\neq\varnothing}{\left(-1+\sum_{\ell=1}^{p}{\textbf{1}[|e\cap P_{\ell}|>0]}\right)}$
$=\sum_{e\in\mathcal{E}^{\prime},e\neq\varnothing}{\left((|e|-1)-\sum_{\ell=1}^{p}{\left(|e\cap P_{\ell}|-\textbf{1}[|e\cap P_{\ell}|>0]\right)}\right)}$
$=\sum_{e\in\mathcal{E}^{\prime},e\neq\varnothing}{\left(\max(|e|-1,0)-\sum_{\ell=1}^{p}{\max(|e\cap P_{\ell}|-1,0)}\right)}$
@@
By Lemma 2.4, there exists a subset $J\subset[L+1]$ of at least two vertices such that $\mathcal{H}^{\prime}=(J,\{J\cap e:e\in\mathcal{E}\})$ — which is exactly the agreement hypergraph of $(y,c^{(j)}:j\in J)$ — is $k$-weakly-partition-connected. ∎
@@ 2.3 Reduced intersection matrices: definition and example
As in [GZ23], we work with the reduced intersection matrix, though our proof should work essentially the same with a different matrix called the (non-reduced) intersection matrix, which was considered in [ST20, ${}^{+}$22, BGM23].
@@
The reduced intersection matrix $\mathsf{RIM}_{k,q,\mathcal{H}}$ associated with a prime power $q$, degree $k$, and a hypergraph $\mathcal{H}=([t],(e_{1},\dots,e_{n}))$ is a $\operatorname*{wt}(\mathcal{E})\times(t-1)k$ matrix over the field of fractions $\mathbb{F}_{q}(X_{1},\dots,X_{n})$. For each hyperedge $e_{i}$ with vertices $j_{1}<j_{2}<\dots<j_{|e_{i}|}$, we add $\operatorname*{wt}(e_{i})=|e_{i}|-1$ rows to $\mathsf{RIM}_{\mathcal{H}}$. For $u=2,\dots,|e_{i}|$, we add a row $r_{i,u}=(r^{(1)},\ldots,r^{(t-1)})$ of length $(t-1)k$ defined as follows:
@@
We typically omit $k$ and $q$ and write $\mathsf{RIM}_{\mathcal{H}}$ as $q$ is typically understood.
@@
$e_{n-2}$$e_{n-1}$$e_{n}$$e_{n-2}$$e_{n-1}$$e_{n}$$f^{(1)}$$f^{(2)}$$f^{(3)}$$f^{(4)}$$f^{(5)}$$f^{(6)}$$f^{(7)}$
The edges $e_{n-2},e_{n-1},e_{n}$ from $\mathcal{H}$ contribute the following length $(t-1)k$ rows to its reduced intersection matrix:
@@
Here $V_{i}=[1,X_{i},X_{i}^{2},\dots,X_{i}^{k-1}]$ is a “Vandermonde row”, and 00 denotes the length-$k$ vector $[0,0,\dots,0]$. Note that each edge $e$ contributes $|e|-1$ rows to the agreement matrix, and in particular $e_{n}$ does not contribute any rows.
Reduced intersection matrices arise by encoding all agreements from a bad list-decoding configuration into linear constraints on the message symbols (the polynomial coefficients). These constraints are placed into one matrix that we call the reduced intersection matrix. The following lemma implies that, if every reduced intersection matrix arising from a possible bad list-decoding configuration has full column rank when $X_{1}=\alpha_{1},\dots,X_{n}=\alpha_{n}$, the corresponding Reed–Solomon code is list-decodable.
@@
Let $\mathcal{H}$ be an agreement hypergraph for $(y,c^{(1)},\dots,c^{(t)})$, where $c^{(j)}\in\mathbb{F}_{q}^{n}$ are codewords of $RS_{n,k}(\alpha_{1},\dots,\alpha_{n})$, not all equal to each other. Then the reduced intersection matrix $\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})$ does not have full column rank.
@@
where $f^{(1)},\dots,f^{(t)}\in\mathbb{F}_{q}^{k}$ are the vectors of coefficients of the polynomials that generate codewords $c^{(1)},\dots,c^{(t)}\in\mathbb{F}_{q}^{n}$. Since these vectors are not all equal to each other, $\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})$ does not have full column rank. ∎
@@
From this definition, it should be clear that we can divide the variables $X_{1},\dots,X_{n}$ into at most $2^{L}$ classes such that variables in the same class are exchangeable with respect to the reduced intersection matrix $\mathsf{RIM}_{\mathcal{H}}$: if $e_{i}$ and $e_{i^{\prime}}$ are the same hyperedge, then swapping $X_{i}$ and $X_{i^{\prime}}$ yields the same reduced intersection matrix (up to row permutations). This observation, which was alluded to in [GZ23], turns out to be crucial in our argument that allows us to improve the alphabet size in [GZ23] from quadratic to linear.

### v6

@@ 2.1 Basic Notation
For positive integers $t$, let $[t]$ denote the set $\{1,2,\dots,t\}$. The Hamming distance $d(x,y)$ between two vectors $x,y\in\mathbb{F}_{q}^{n}$ is the number of indices $i$ where $x_{i}\neq y_{i}$. For a finite field $\mathbb{F}_{q}$, we follow the standard notation that $\mathbb{F}_{q}[X_{1},\dots,X_{n}]$ denotes the ring of multivariate polynomials with variables $X_{1},\dots,X_{n}$ over $\mathbb{F}_{q}$, and $\mathbb{F}_{q}(X_{1},\dots,X_{n})$ denotes the field of fractions of the polynomial ring $\mathbb{F}_{q}[X_{1},\dots,X_{n}]$. By abuse of notation, we let $X_{\leq i}$ or $X_{[i]}$ to denote the sequence $X_{1},\dots,X_{i}$, and we let, for example, $X_{\leq i}=\alpha_{\leq i}$ to denote $X_{1}=\alpha_{1},X_{2}=\alpha_{2},\dots,X_{i}=\alpha_{i}$. Given a matrix $M$ over the field of fractions $\mathbb{F}_{q}(X_{1},\dots,X_{n})$ whose entries are in $\mathbb{F}_{q}[X_{1},\dots,X_{n}]$ and field elements $\alpha_{1},\dots,\alpha_{i}\in\mathbb{F}_{q}$, let $M(X_{\leq i}=\alpha_{\leq i})$ denote the matrix over $\mathbb{F}_{q}(X_{i+1},X_{i+2},\dots,X_{n})$ obtained by setting $X_{\leq i}=\alpha_{\leq i}$ in $M$.
@@ 2.2 Hypergraphs and Connectivity
In this work, we maintain a hypergraph perspective of the list-decoding problem, which was introduced in [25]. We describe a bad list-decoding instance with a hypergraph where the $L+1$ bad codewords identify the vertices and the $n$ evaluation points identify the hyperedges (Definition 2.1). While prior works described a bad list-decoding instance by $L+1$ sets indicating the agreements of the codewords with the received word, this hypergraph perspective gives us several advantages:
@@
    The constraints imposed by a bad list-decoding configuration yield a hypergraph that is weakly-partition-connected. This is a natural notion of hypergraph connectivity, which is well-studied in combinatorics [23, 22, 45] and optimization [43, 21, 20, 11], and which generalizes a well-known notion ($k$-partition-connectivity) for graphs [52, 62]. [^6] The notion of weakly-partition-connected sits between two other well-studied notions: $k$-partition-connected implies $k$-weakly-partition-connected implies $k$-edge-connected [45]. Each of these three notions generalizes an analogous notion on graphs. On graphs, $k$-partition-connected and $k$-weakly-partition-connected are equivalent. This connection allows us to have more compact notation, definitions, and lemma statements, and allows us to streamline some proofs.
@@
    With the hypergraph perspective, we can give a new presentation of the results in [9] and more directly connect the list-decoding problem to the GM-MDS theorem [13, 49, 65], as the heavy-lifting in the combinatorics is done using known results on hypergraph orientations. This is done in Appendix A.
A hypergraph $\mathcal{H}=(V,\mathcal{E})$ is given by a set of vertices $V$ and a set $\mathcal{E}$ of (hyper)edges, which are (possibly) subsets of the vertices $V$. In this work, all hypergraphs have labeled edges, meaning we enumerate our edges $e_{i}$ by distinct indices $i$ from some set, typically $[n]$, in which case we may also think of $\mathcal{E}$ as a tuple $(e_{1},\dots,e_{n})$. Throughout this paper, the vertex set $V$ is typically $[t]$ for some positive integer $t$. The weight of a hyperedge $e$ is $\operatorname*{wt}(e)\stackrel{{\scriptstyle\rm def}}{{=}}\max(0,|e|-1)$, and the weight of a set of hyperedges $\mathcal{E}$ is simply $\operatorname*{wt}(\mathcal{E})\stackrel{{\scriptstyle\rm def}}{{=}}\sum_{e\in\mathcal{E}}{\operatorname*{wt}(e)}$.
$e_{n-2}$$e_{n-1}$$e_{n}$$f^{(1)}$$f^{(2)}$$f^{(3)}$$f^{(4)}$$f^{(5)}$$f^{(6)}$$f^{(7)}$$e_{n-2}=\{1,2,4\}$ means $f^{(1)}(\alpha_{n-2})=f^{(2)}(\alpha_{n-2})=f^{(4)}(\alpha_{n-2})=y_{n-2}$$e_{n-1}=\{5,6\}$ means $f^{(5)}(\alpha_{n-1})=f^{(6)}(\alpha_{n-1})=y_{n-1}$$e_{n}=\{7\}$ means $f^{(7)}(\alpha_{n})=y_{n}$
@@
Agreement hypergraphs (or their subgraphs) enjoy a key property called weak-partition-connectivity.
Definition 2.2 (Weak-Partition-Connectivity).
@@
where $|\mathcal{P}|$ is the number of parts of the partition, and $|\mathcal{P}(e)|$ is the number of parts of the partition that edge $e$ intersects. [^7] We take $\max\{|\mathcal{P}(e)|-1,0\}$ rather than $|\mathcal{P}(e)|-1$ because the hyperedges in agreement hypergraphs could be empty.
To give some intuition for weak-partition-connectivity, we state two of its combinatorial implications. First, if a hypergraph is $k$-weakly-partition-connected, then it is $k$-edge-connected [45], which, by the Hypergraph Menger’s (Max-Flow-Min-Cut) theorem [45, Theorem 1.11], equivalently means that every pair of vertices has $k$ edge-disjoint (hyper)paths between them. [^8] In general the converse is not true. Second, suppose we replace every hyperedge $e$ with an arbitrary spanning tree of its vertices (which we effectively do in Definition 2.6). The resulting (non-hyper)graph is $k$-partition-connected, [^9] In (non-hyper)graphs, $k$-partition-connectivity and $k$-weak-partition-connectivity are equivalent. which, by the Nash-Williams-Tutte Tree-Packing theorem [52, 62], equivalently means there are $k$ edge-disjoint spanning trees (this connection was used in [25]).
@@
Lemma 2.3 (Bad list gives $k$-weakly-partition-connected hypergraph. See also [25, Lemma 7.4]).
@@
Lemma 2.3 follows from the following result about weakly-partition-connected hypergraphs
@@
Let $\mathcal{H}=(V,\mathcal{E})$ be a hypergraph with at least two vertices such that $\sum_{e\in\mathcal{E}}\operatorname*{wt}(e)\geq k\cdot(|V|-1)$, where $k$ is a positive integer. Then there exists a subset $V^{\prime}\subseteq V$ of at least two vertices such that the hypergraph $\mathcal{H}^{\prime}=(V^{\prime},\{e\cap V^{\prime}:e\in\mathcal{E}\})$ is $k$-weakly-partition-connected.
@@
By assumption, $V^{\prime}=[L+1]$ satisfies (3), so $V^{\prime}$ exists (note that singleton subsets of $[L+1]$ satisfy (3) with equality). Let $\mathcal{H}=(V^{\prime},\mathcal{E}^{\prime})$ be the hypergraph with edge set $\mathcal{E}^{\prime}=\{V^{\prime}\cap e:e\in\mathcal{E}\}$. By minimality of $V^{\prime}$, for all nonempty $V^{\prime\prime}\subsetneq V^{\prime}$, we have $\sum_{e\in\mathcal{E}^{\prime}}{\operatorname*{wt}(e\cap V^{\prime\prime})}\leq(|V^{\prime\prime}|-1)k$. Now, consider a non-trivial partition $\mathcal{P}=P_{1}\sqcup\cdots\sqcup P_{p}$ of $V^{\prime}$ where $P_{i}\neq V^{\prime}$ for all $i\in[p]$ (as otherwise (2) trivially follows). We have
$\sum_{e\in\mathcal{E}^{\prime}}{\max\{|\mathcal{P}(e)|-1,0\}}$ $=\sum_{e\in\mathcal{E}^{\prime}}\left(\operatorname*{wt}(e)-\sum_{\ell=1}^{p}{\operatorname*{wt}(e\cap P_{\ell})}\right)$
@@
By Lemma 2.4, there exists a subset $J\subseteq[L+1]$ of at least two vertices such that $\mathcal{H}^{\prime}=(J,\{J\cap e:e\in\mathcal{E}\})$ — which is exactly the agreement hypergraph of $(y,c^{(j)}:j\in J)$ — is $k$-weakly-partition-connected. ∎
@@ 2.3 Reduced Intersection Matrices: Definition and Example
We work with the reduced intersection matrix, which encodes all agreements from a bad list-decoding configuration into linear constraints on the message symbols (the polynomial coefficients). We point out that previous works [59, 25, 9] considered a related matrix called the (non-reduced) intersection matrix, and our proof would work just as well with this matrix.
@@
The reduced intersection matrix $\mathsf{RIM}_{k,q,\mathcal{H}}$ associated with a prime power $q$, degree $k$, and a hypergraph $\mathcal{H}=([t],(e_{1},\dots,e_{n}))$ is a $\operatorname*{wt}(\mathcal{E})\times(t-1)k$ matrix over the field of fractions $\mathbb{F}_{q}(X_{1},\dots,X_{n})$. It is constructed as follows. For each hyperedge $e_{i}$ with vertices $j_{1}<j_{2}<\dots<j_{|e_{i}|}$, we add $\operatorname*{wt}(e_{i})=|e_{i}|-1$ rows to $\mathsf{RIM}_{k,q,\mathcal{H}}$. For $u=2,\dots,|e_{i}|$, we add a row $r_{i,u}=(r^{(1)},\ldots,r^{(t-1)})$ of length (t1)k(t-1)k defined as follows:
@@
We typically omit $k$ and $q$ and write $\mathsf{RIM}_{\mathcal{H}}$ as $k$ and $q$ are typically understood.
@@
$e_{n-2}$$e_{n-1}$$e_{n}$$f^{(1)}$$f^{(2)}$$f^{(3)}$$f^{(4)}$$f^{(5)}$$f^{(6)}$$f^{(7)}$
The edges $e_{n-2},e_{n-1},e_{n}$ from $\mathcal{H}$ contribute the following length (t1)k(t-1)k rows to its reduced intersection matrix:
@@
Here $V_{i}=[1,X_{i},X_{i}^{2},\dots,X_{i}^{k-1}]$ is a “Vandermonde row”, and 0 denotes the length-$k$ vector $0,0,\dots,0]$. Note that each edge $e$ contributes |e|1|e|-1 rows to the agreement matrix, and in particular $e_{n}$ does not contribute any rows.
The following lemma implies that, if every reduced intersection matrix arising from a possible bad list-decoding configuration has full column rank when $X_{1}=\alpha_{1},\dots,X_{n}=\alpha_{n}$, the corresponding Reed–Solomon code is list-decodable.
@@
Let $\mathcal{H}$ be an agreement hypergraph for $(y,c^{(1)},\dots,c^{(t)})$, where $c^{(j)}\in\mathbb{F}_{q}^{n}$ are codewords of $\mathsf{RS}_{n,k}(\alpha_{1},\dots,\alpha_{n})$, not all equal to each other. Then the reduced intersection matrix $\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})$ does not have full column rank.
@@
where $f^{(1)},\dots,f^{(t)}\in\mathbb{F}_{q}^{k}$ are the vectors of coefficients of the polynomials that generate the codewords $c^{(1)},\dots,c^{(t)}\in\mathbb{F}_{q}^{n}$. Since these vectors are not all equal to each other, $\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})$ does not have full column rank. ∎
@@
From this definition, it should be clear that we can divide the variables $X_{1},\dots,X_{n}$ into at most $2^{L}$ classes such that variables in the same class are exchangeable with respect to the reduced intersection matrix $\mathsf{RIM}_{\mathcal{H}}$: if $e_{i}$ and $e_{i^{\prime}}$ are the same hyperedge, then swapping $X_{i}$ and $X_{i^{\prime}}$ yields the same reduced intersection matrix (up to row permutations). This observation turns out to be crucial for bringing the alphabet size all the way down to linear; without it, we would get a quadratic alphabet size.